At the end of its first month of operations, JMP Consulting reported net income of $29,500. They also had account balances of: Cash, $21,000; Office Supplies, $2,750 and Accounts Receivable, $11,500. The owner's total investment for this first month was $5,750. There were no owner withdrawals in the first month.

Calculate the ending balance in the Owner's Capital account to be reported on the Statement of Owner's Equity.

2. A company purchased a new delivery van at a cost of $47,000 on July 1. The delivery van is estimated to have a useful life of 5 years and a salvage value of $3,500. The company uses the straight-line method of depreciation. How much depreciation expense will be recorded for the van during the first year ended December 31?

2.  
Answer #1

1) Total assets on First month = 21000+2750+11500 = 35250

Total Ending balance of owner's capital :

Beginning balance 5750
Add: Net income 29500
Less: Withdrawal 0
Ending balance 35250

2) depreciation expense for first year = (47000-3500)/5*6/12 = 4350